celticcat:

If x and y are both negative show that x/y + y/x > 2

Zepdrix:

\[\large\rm \frac xy+\frac yx>2\]Multiply both sides by x, (Since x is negative, the inequality flips), \[\large\rm \frac{x^2}y + y<2x\]Multiply both sides by y, (y is negative, inequality flips again), \[\large\rm x^2+y^2>2xy\]Subtract,\[\large\rm x^2-2xy+y^2>0\]These terms can factor down into a perfect square, \[\large\rm (x-y)^2>0\]And since a squared term is always positive (always greater than zero), this holds true. `If x and y are both negative show that x/y + y/x > 2` To properly "show" that, we would want to apply all of these steps in reverse. We would start with (x-y)^2>0. Mention that it holds true because it's a square. And show that it expands to the equation we're looking for.

celticcat:

THanks a lot . I always have a hard time with this type of question.

sillybilly123:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @celticcat If x and y are both negative show that x/y + y/x > 2 \(\color{#0cbb34}{\text{End of Quote}}\) A Contradiction: let \(x = y = \alpha, \qquad \alpha \ne 0\) \( \implies \dfrac xy+\dfrac yx = 2 \implies \dfrac xy+\dfrac yx \ngtr 2\) So, Poof(!), it's wrong - this is math. And there are an infinite number of ways that the original proposition fails :( There is also a singularity at the Origin IMHO: using algebra here, is fine but hardly inspired. And using kids' rules, about flipping the sign of the inequality is, ..... mmmm, interesting. not that interested in doing so, but i reckon you could work the proof in one original direction without the a priori back-tracking that the **5-star** answer suggests 😂 occurred to me that you could write it as \(u + \dfrac{1}{u} \gt 2\), then plot it as a single variable problem: that is an interesting approach, maybe